# Automated Black-box Prompt Engineering for Personalized Text-to-Image Generation

### v3

@@ 1 Introduction
An important goal of generative modeling is to design algorithms capable of steering generative models to produce desired output images. Early attempts, which often centered on particular architectures or tasks, were largely characterized by manually-curated data collection, fine-tuning, or retraining from scratch (Srivastava & Salakhutdinov, 2012; Mirza & Osindero, 2014; Isola et al., 2017; Zhu et al., 2017). These requirements are often costly, and the resulting solutions usually do not transfer well between models. Thus despite the promise of these methods, efficient and generalized algorithms for controllable generation remain sought after.
@@ 3.2 Algorithm
1:Input: $N$ streams, $K$ iterations, $\{x_{i}\}_{i=1}^{M}$ reference images
@@ 4.4 Ablation Study
Figure 7: Ablation study on different numbers of streams $N$ with fixed $K=5$.
@@
Next we take a closer look at the effect of increasing the budget in PRISM. Figure 8 and 8 show the effect of increasing the number of streams $N$ and the number of iterations $K$ respectively. We observe that when increasing $N$ and keeping $K$ fixed, we can obtain steady performance improvements in both human readability and prompt accuracy. When increasing $K$ and keeping $N$ fixed, although we do not observe a monotonic relationship between the performance and $K$, we can still notice a general upward trend in prompt accuracy. In Appendix D, we discuss the trade-off between $N$ and $K$ to better inform practitioners how to choose these hyperparameters. Besides adjusting the budget, one can also use cheaper or open-sourced VLMs to lower the cost. In Appendix C.2, we experiment with GPT-4o-mini and IDEFICS2 (Laurençon et al., 2024), two significantly cheaper and smaller VLMs to demonstrate the cost-flexibility of PRISM.
@@ 4.5 Additional Applications
Unlike PEZ, which requires an additional optimization process to generate distilled prompts, PRISM leverages its highly interpretable prompts and VLMs to simplify prompts effectively. By directly instructing a VLM (in this case GPT-4o) with the prompt “Here is a prompt to this image with a text-to-image model, make it more concise (less than $<$length$>$ tokens) but keep all the descriptive details”, we achieve concise, distilled prompts without additional computational overhead, as shown in Figure 11.
@@ Appendix A Additional Experiment Details
To simplify the implementation, we only keep a chat history length of $3$ and use the length of the prompt as an approximation of the log-likelihood for the final prompt selection. When evaluating the judge scores $\mathbf{D}(x,\hat{x})$ in PRISM iterations, we shuffle the reference images when $M>1$. The judge score is rescaled into a range from 0 to 10. For direct image inversion, we re-evaluate the top $5$ candidates twice and tally the scores with in-iteration scores to make the final decision. For personalized T2I generation, we re-evaluate once for each reference image and use the average score to select the output.
@@ Appendix B Designing System Prompts
System prompting is the standard way to condition a general purpose LLM for specific tasks of request. The key idea is that, before the conversation starts, the LLM receives a tailored message, the system prompt, that provides the contexts, conversation scenario settings, formats and other guidelines as the prefix of the entire conversation ahead. In this section, we elaborate on the design of the system prompts for the prompt engineer assistant ${\mathbf{F}}$ and the judge ${\mathbf{D}}$. We also provide the full system prompts used in all of our experiments at the end of this paper in Section G and in our demo code.
@@ Appendix E Limitations and Future Works
Finally, we want to re-iterate the potential societal impacts of our work. Just like LLMs are prone to jail-breaking and leaking, we also do not guarantee complete protection against malicious use intent, underlying bias and other limitations inherent from the base models. We are committed to implement and constantly improve the safety precautions in our code base after its public release, and we encourage practitioners to also take preventative actions in order to mitigate these potential issues.

### v4

@@ 1 Introduction
An important goal of generative modeling is to design algorithms capable of steering generative models to produce desired output images. Early attempts, which often centered on particular architectures or tasks, were largely characterized by manually-curated data collection, fine-tuning, or retraining from scratch (Srivastava & Salakhutdinov, 2012; Mirza & Osindero, 2014; Isola et al., 2017; Zhu et al., 2017). These requirements are often costly, and the solutions usually do not transfer well between models. Thus despite the promise of these methods, efficient and generalized algorithms for controllable generation remain sought after.
@@ 3.2 Algorithm
1:Input: NN streams, $K$ iterations, $\{x_{i}\}_{i=1}^{M}$ reference images
@@ 4.4 Ablation Study
Figure 7: Ablation study on different numbers of streams NN with fixed $K=5$.
@@
Next we take a closer look at the effect of increasing the budget in PRISM. Figure 8 and 8 show the effect of increasing the number of streams $N$ and the number of iterations $K$ respectively. We observe that when increasing $N$ and keeping $K$ fixed, we can obtain steady performance improvements in both human readability and prompt accuracy. When increasing $K$ and keeping $N$ fixed, although we do not observe a monotonic relationship between the performance and $K$, we can still notice a general upward trend in prompt accuracy. In Appendix D, we discuss the trade-off between $N$ and $K$ to better inform practitioners how to choose these hyperparameters. Besides adjusting the budget, one can also use cheaper or open-sourced VLMs to lower the cost. A typical PRISM run with GPT-4V and $N\times K=30$ budget costs around \$1.5 USD and 11 minutes (as of May 2024). However, in Appendix C.2, we demonstrate that PRISM can still achieve high performance with GPT-4o-mini ($2\%$ of the price of GPT-4V) and IDEFICS2 (Laurençon et al., 2024) (free of charge), showing the cost-flexibility of our method. In Appendix E, we provide detailed comparison regarding cost and latency among baselines and PRISM with different VLMs.
@@ 4.5 Additional Applications
Unlike PEZ, which requires an additional optimization process to generate distilled prompts, PRISM leverages its highly interpretable prompts and VLMs to simplify prompts effectively. By directly instructing a VLM (in this case GPT-4o) with the prompt “Here is a prompt to this image with a text-to-image model, make it more concise (less than <<length>> tokens) but keep all the descriptive details”, we achieve concise, distilled prompts without additional computational overhead, as shown in Figure 11.
@@ Appendix A Additional Experiment Details
[…]
@@ Appendix B Designing System Prompts
System prompting is the standard way to condition a general purpose LLM for specific tasks of request. The key idea is that, before the conversation starts, the LLM receives a tailored message, the system prompt, that provides the contexts, conversation scenario settings, formats and other guidelines as the prefix of the entire conversation ahead. In this section, we elaborate on the design of the system prompts for the prompt engineer assistant $\mathbf{F}$} and the judge $\mathbf{D}$}. We also provide the full system prompts used in all of our experiments at the end of this paper in Section G.
@@ C.2 Flexible Model Choices
C.3 Discussions on prompts rejected by T2I safeguards
Moreover, as previously demonstrated, while our method is generally safer than most baselines, there are still instances where the generated prompts fail to bypass the safeguards of the T2I model. For example, the prompt “single three-dimensional emoji-like object, stylized pile of poop, very smooth and uniform matte surface, solid light brown color without any pattern, cartoon eyes accurately sized, simple mouth with a cheerful expression, isolated” is rejected by DALL-E 2’s content filters. This failure is likely due to the presence of certain keywords (e.g., “poop”) that trigger automated content moderation, regardless of the otherwise harmless intent of the prompt.
On the contrary, baseline methods like CLIP-Interrogator often generate prompts that include the names of specific artists or copyrighted entities. This issue arises from their use of a pre-collected keyword bank that contains numerous proper nouns, including well-known artist names and intellectual property (IP) references like “Pokemon”. Prompts generated in this way are not only more likely to be rejected by models such as DALL-E, which explicitly ban names of living artist and copyrighted IP, but also carry a higher risk of producing outputs that closely mimic copyrighted styles or identities. This substantially increases the potential for copyright infringement, making these methods less suitable for deployment in practical settings where safety and legal compliance are critical.
@@ Appendix E Limitations and Future Works
As we have observed in Section D, different choices of the base model significantly affects the performance of PRISM. For example, when using CLIP as the Judge model, our method can fail to capture fine-grained details or distinguish between the main object and the irrelevant background due to the limitation of the CLIP model. Moreover, when using a significantly smaller VLM such as IDEFICS2, the performance can also be slightly degraded. Overall, one should stay cautious when choosing the base models for our PRISM, yet as we have mentioned above, cheaper models such as GPT-4o-mini is already sufficient for achieving comparable results, and we remain hopeful for small open-sourced model to catch up with GPT-4V in the future.
Finally, we want to re-iterate the potential societal impacts of our work. Just like LLMs are prone to jail-breaking and leaking, we also do not guarantee complete protection against malicious use intent, underlying bias and other limitations inherent from the base models. For example, common jail-breaking techniques such as GCG (Zou et al., 2023) and PAIR (Chao et al., 2023) can be used to find harmful prompt to elicit contents with nudity or gore. These jail-breaking methods can also be applied in the Judge model to enforce certain preferences with mal-intent. We are committed to implement and constantly improve the safety precautions in our code base after its public release, and we encourage practitioners to also take preventative actions in order to mitigate these potential issues.